OpenStudy (anonymous):

An inlet pipe can fill an empty swimming pool in 7 hours, and another inlet pipe can fill the pool in 2 hours. How long will it take both pipes to fill the pool

OpenStudy (anonymous):

1/7 + 1/2 = 1/x after solving x = 1.55 or 1 hr 33 minutes

OpenStudy (anonymous):

thats what I got but my online retriceignment says its wrong

OpenStudy (anonymous):

pls check wheteher it is ...........one pipe is inlet n other for outlet

OpenStudy (anonymous):

i dunno know thanks anyway

OpenStudy (anonymous):

if it so then try..1/7 - 1/2 = 1/x...it should work out